# Extending Graph Condensation to Multi-Label Datasets: A Benchmark Study

### v1

@@ 5 Datasets and Experiment Settings
  • PPI Zeng et al. (2019) is the Protein-Protein Interaction network. It classifies protein functions based on the interactions of human tissue proteins. Positional gene sets are used, motif gene sets and immunological signatures as features and gene ontology sets as labels (121 in total), collected from the Molecular Signatures Database Liberzon et al. (2015). And PPI-large is the larger version of PPI dataset.
@@
  • Yelp Zeng et al. (2019) is processed from a public open dataset from yelp.com. The nodes represent active users and the edges are the relationship between the users. The multi-label of each node represents the types of business like Coffee & Tea, Flowers & Gifts, Tours, and so on. Node features contain the information of all the reviews and rates by users, generated by pre-trained Word2Vec model Church (2017).
@@
  • DBLP Akujuobi et al. (2019) is the citation network extracted from DBLP. The nodes represent authors and edges are the co-authorship between the authors. Multi-label here indicates the four research areas like database, data mining, information and retrieval and artificial intelligence.
@@
  • OGBN-Proteins from Open Graph Benchmark Hu et al. (2020) is an undirected, weighted, and typed (according to species) graph. Nodes represent proteins, and edges indicate different types of biologically meaningful associations between proteins, e.g., physical interactions, co-expression or homology. The task is to predict the presence of protein functions in a multi-label binary classification setup, where there are 112 kinds of protein functions denote as labels to predict in total. Here the protein functions refer to the specific biological activities or roles that a protein performs within a cell or organism. These functions are typically described using Gene Ontology (GO) database Huntley et al. (2015) or other biological annotation systems and can fall into different categories.
@@
  • PCG Zhao et al. (2023) is Protein-Phenotype graph dataset, which focused on predicting phenotypes associated with proteins. Phenotypes refer to observable traits or characteristics of diseases, and identifying these associations can be valuable for clinical diagnostics or discovering potential drug targets. The dataset is constructed using protein-phenotype associations from the DisGeNET database Piñero et al. (2020), where each protein is linked to one or more phenotypes. The phenotypes are grouped into disease categories based on the MESH ontology Bhattacharya et al. (2011), and labels with fewer than 100 associated proteins are removed. The PCG includes 3,233 proteins as nodes and 37,351 protein-protein interactions as edges. The task is to predict multiple phenotype labels for each protein based on its interactions and features.
@@
  • HumanGo Chou & Shen (2007); Liberzon et al. (2015) contains 3,106 proteins, each potentially associated with one or more of 14 subcellular locations as multi-label. The proteins are represented as nodes in a graph, with features generated from their sequences using pre-trained model Yang et al. (2020). Protein-protein interactions form the edges of the graph, with each edge representing the confidence of different interaction types. Different location associations come from the GO database Huntley et al. (2015), which assigns standardized terms to proteins based on their roles in biological processes (e.g., cell division, metabolism), molecular functions (e.g., enzyme activity, receptor binding), and cellular components (e.g., nucleus, membrane).
@@
  • EukaryoteGo Chou & Shen (2007); Liberzon et al. (2015) follows a similar structure but focuses on eukaryote proteins. It includes 7,766 proteins and their interactions, with proteins being assigned to one or more of 22 subcellular locations. Like HumanGo, it uses protein sequences for node features and protein-protein interactions for graph structure.
@@
We follow the predefined data splits from Zeng et al. (2019); Zhao & Bilen (2023); Hu et al. (2020); Chou & Shen (2007). 1 presents an overview of the datasets’ characteristics.
@@
[…]
@@
[…]
@@ 6.1 Multi-label Adaptation
As some of the coreset methods gain comparable performance as baseline Welling (2009); Farahani & Hekmatfar (2009); Sener & Savarese (2017), we apply those heuristic algorithms to extract $N^{\prime}$ nodes from the full set of nodes to obtain subgraphs $S^{\prime}$ with corresponding condensation rates as the initialization methods.
@@
  • Herding Welling (2009) method is more structured compared with random selection strategy. It is based on iterative minimization of the distance between the mean of the selected nodes’ feature vectors and the mean of the entire graph feature vectors. Let $\mu_{\mathcal{G}}$ present the mean feature vector of the original graph:
@@
  • K-Center Farahani & Hekmatfar (2009); Sener & Savarese (2017) method aims to maximize the coverage of the original graph feature space by minimizing the maximum distance between unselected nodes and their closest selected node. Let $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{N^{\prime}}\}$ be the selected node set. The goal is to select nodes such that for every unselected node $i$, the distance to the nearest selected node $s_{j}\in\mathcal{S}$ is minimized:
@@
Above all, in multi-label settings, we adapt subgraph initialization and probabilistic synthetic multi-label methods for multi-label graph condensation. Besides, in graph condensation, the structure $A^{\prime}$ of synthetic graph is optimized from nodes features $X^{\prime}$. Based on the researches Ding et al. (2022); Zhu et al. (2021); Jin et al. (2022b), this "Graphless" learning process can be highly useful for downstream data analysis, which also shows competitive performance. Therefore, by comparing different initialization methods with or without learning the graph structure, we get the best settings for the adaptation.
@@
However, in multi-label classification $Y^{\prime}\in\{y_{1}^{\prime},\cdots,y_{N}^{\prime}\}\in\{0,1\}^{N^{\prime}\times K}$, each node can associate with multiple classes simultaneously. Typically, the multi-label classification task is modeled as $K$ independent binary classification tasks, for each task $k$, the model predicts a logit $z_{j,k}$ and ground-truth label $y_{j,k}$ indicates whether node $j$ belongs to class $k$. In multi-label classification, SoftMarginLoss Cao et al. (2019) incorporates margin-based optimization, which is beneficial when dealing with imbalanced datasets and overlapping classes. In multi-label settings, certain labels may frequently co-occur or exhibit correlations. SoftMarginLoss introduces a margin between positive and negative label predictions, encouraging the model to differentiate between these labels more confidently.
@@
This formulation is particularly useful in ensuring that the model not only predicts correct labels but does so with confidence, reducing the chances of ambiguous predictions near the decision boundary. On the other hand, Binary Cross-Entropy Loss Durand et al. (2019) is a widely-used objective function for multi-label classification, as it independently evaluates each label for each node in the graph. Then the task-wise loss for each binary classification task is computed using Binary Cross-Entropy (BCE) Durand et al. (2019) loss:
@@ 6.2 Experiments and Results Analysis
Specifically, we report the performance of various initialization strategies include subgraph sampling( Random, Herding, K-Center) Welling (2009); Farahani & Hekmatfar (2009); Sener & Savarese (2017), and probabilistic label sampling for synthetic $Y^{\prime}$. After the initialization, the synthetic graph $\mathcal{S}$ will be optimized by gradient matching strategy in GCond. Then we compare different optimization loss functions, including SoftMarginLoss Cao et al. (2019) and BCE loss Durand et al. (2019), across diverse datasets with pre-choosen condensation rates. Through these experiments, we aim to identify the optimal combination of initialization and loss function that achieves the best trade-off between graph size reduction and model performance in multi-label tasks.

### v2

@@ 5 Datasets and Experiment Settings
  • PPI zeng2019graphsaint is the Protein-Protein Interaction network. It classifies protein functions based on the interactions of human tissue proteins. Positional gene sets are used, motif gene sets and immunological signatures as features and gene ontology sets as labels (121 in total), collected from the Molecular Signatures Database liberzon2015molecular. And PPI-large is the larger version of PPI dataset.
@@
  • Yelp zeng2019graphsaint is processed from a public open dataset from yelp.com. The nodes represent active users and the edges are the relationship between the users. The multi-label of each node represents the types of business like Coffee & Tea, Flowers & Gifts, Tours, and so on. Node features contain the information of all the reviews and rates by users, generated by pre-trained Word2Vec model church2017word2vec.
@@
  • DBLP akujuobi2019collaborative is the citation network extracted from DBLP. The nodes represent authors and edges are the co-authorship between the authors. Multi-label here indicates the four research areas like database, data mining, information and retrieval and artificial intelligence.
@@
  • OGBN-Proteins from Open Graph Benchmark hu2020ogb is an undirected, weighted, and typed (according to species) graph. Nodes represent proteins, and edges indicate different types of biologically meaningful associations between proteins, e.g., physical interactions, co-expression or homology. The task is to predict the presence of protein functions in a multi-label binary classification setup, where there are 112 kinds of protein functions denote as labels to predict in total. Here the protein functions refer to the specific biological activities or roles that a protein performs within a cell or organism. These functions are typically described using Gene Ontology (GO) database huntley2015goa or other biological annotation systems and can fall into different categories.
@@
  • PCG zhao2023multi is Protein-Phenotype graph dataset, which focused on predicting phenotypes associated with proteins. Phenotypes refer to observable traits or characteristics of diseases, and identifying these associations can be valuable for clinical diagnostics or discovering potential drug targets. The dataset is constructed using protein-phenotype associations from the DisGeNET database pinero2020disgenet, where each protein is linked to one or more phenotypes. The phenotypes are grouped into disease categories based on the MESH ontology bhattacharya2011mesh, and labels with fewer than 100 associated proteins are removed. The PCG includes 3,233 proteins as nodes and 37,351 protein-protein interactions as edges. The task is to predict multiple phenotype labels for each protein based on its interactions and features.
@@
  • HumanGo chou2007euk; liberzon2015molecular contains 3,106 proteins, each potentially associated with one or more of 14 subcellular locations as multi-label. The proteins are represented as nodes in a graph, with features generated from their sequences using pre-trained model yang2020prediction. Protein-protein interactions form the edges of the graph, with each edge representing the confidence of different interaction types. Different location associations come from the GO database huntley2015goa, which assigns standardized terms to proteins based on their roles in biological processes (e.g., cell division, metabolism), molecular functions (e.g., enzyme activity, receptor binding), and cellular components (e.g., nucleus, membrane).
@@
  • EukaryoteGo chou2007euk; liberzon2015molecular follows a similar structure but focuses on eukaryote proteins. It includes 7,766 proteins and their interactions, with proteins being assigned to one or more of 22 subcellular locations. Like HumanGo, it uses protein sequences for node features and protein-protein interactions for graph structure.
@@
We follow the predefined data splits from zeng2019graphsaint; zhao2023dataset; hu2020ogb; chou2007euk. 1 presents an overview of the datasets’ characteristics.
@@
    $\text{F1-micro}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (14)
@@
    $\text{F1-macro}=\frac{1}{C}\sum_{i=1}^{C}\text{F1}_{i}$ (15)
    where $C$ is the number of classes and $\text{F1}_{i}$ is the F1 score for the $i^{th}$ class, calculated as:
    $\text{F1}_{i}=\frac{2\times\text{Precision}_{i}\times\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (16)
@@ 6.1 Multi-label Adaptation
As some of the coreset methods gain comparable performance as baseline welling2009herding; farahani2009facility; sener2017active, we apply those heuristic algorithms to extract $N^{\prime}$ nodes from the full set of nodes to obtain subgraphs $S^{\prime}$ with corresponding condensation rates as the initialization methods.
@@
  • Herding welling2009herding method is more structured compared with random selection strategy. It is based on iterative minimization of the distance between the mean of the selected nodes’ feature vectors and the mean of the entire graph feature vectors. Let $\mu_{\mathcal{G}}$ present the mean feature vector of the original graph:
@@
  • K-Center farahani2009facility; sener2017active method aims to maximize the coverage of the original graph feature space by minimizing the maximum distance between unselected nodes and their closest selected node. Let $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{N^{\prime}}\}$ be the selected node set. The goal is to select nodes such that for every unselected node $i$, the distance to the nearest selected node $s_{j}\in\mathcal{S}$ is minimized:
@@
Above all, in multi-label settings, we adapt subgraph initialization and probabilistic synthetic multi-label methods for multi-label graph condensation. Besides, in graph condensation, the structure $A^{\prime}$ of synthetic graph is optimized from nodes features $X^{\prime}$. Based on the researches ding2022data; zhu2021survey; GCond, this "Graphless" learning process can be highly useful for downstream data analysis, which also shows competitive performance. Therefore, by comparing different initialization methods with or without learning the graph structure, we get the best settings for the adaptation.
@@
However, in multi-label classification $Y^{\prime}\in\{y_{1}^{\prime},\cdots,y_{N}^{\prime}\}\in\{0,1\}^{N^{\prime}\times K}$, each node can associate with multiple classes simultaneously. Typically, the multi-label classification task is modeled as $K$ independent binary classification tasks, for each task $k$, the model predicts a logit $z_{j,k}$ and ground-truth label $y_{j,k}$ indicates whether node $j$ belongs to class $k$. In multi-label classification, SoftMarginLoss cao2019learning incorporates margin-based optimization, which is beneficial when dealing with imbalanced datasets and overlapping classes. In multi-label settings, certain labels may frequently co-occur or exhibit correlations. SoftMarginLoss introduces a margin between positive and negative label predictions, encouraging the model to differentiate between these labels more confidently.
@@
This formulation is particularly useful in ensuring that the model not only predicts correct labels but does so with confidence, reducing the chances of ambiguous predictions near the decision boundary. On the other hand, Binary Cross-Entropy Loss durand2019learning is a widely-used objective function for multi-label classification, as it independently evaluates each label for each node in the graph. Then the task-wise loss for each binary classification task is computed using Binary Cross-Entropy (BCE) durand2019learning loss:
@@ 6.2 Experiments and Results Analysis
Specifically, we report the performance of various initialization strategies include subgraph sampling( Random, Herding, K-Center) welling2009herding; farahani2009facility; sener2017active, and probabilistic label sampling for synthetic $Y^{\prime}$. After the initialization, the synthetic graph $\mathcal{S}$ will be optimized by gradient matching strategy in GCond. Then we compare different optimization loss functions, including SoftMarginLoss cao2019learning and BCE loss durand2019learning, across diverse datasets with pre-choosen condensation rates. Through these experiments, we aim to identify the optimal combination of initialization and loss function that achieves the best trade-off between graph size reduction and model performance in multi-label tasks.